OpenStudy (anonymous):

Evaluate f(x) for the value indicated. f(x) = -2x^2 - 9x + 4 for f(-2)

OpenStudy (anonymous):

when you have f(-2) it simply means that solve f(x) or y when x=-2

OpenStudy (anonymous):

should get 5 from just doing it in myhead

OpenStudy (anonymous):

14, 18, -22, or 30. I am going with 14.

OpenStudy (anonymous):

wait are you sure you have the right equation up there

OpenStudy (anonymous):

-8+18+4

OpenStudy (anonymous):

yes it's 14 my bad forgot to times by

OpenStudy (anonymous):

2

OpenStudy (anonymous):

This is what she gave me.

OpenStudy (anonymous):

think it's bed time

OpenStudy (anonymous):

f(x) = -2x^2 - 9x + 4 for f(-2) f(-2)=-2(-2)^2-9(-2)=-8+18=10